OpenStudy (anonymous):

Write down an equation of circle which is concentric with the circle x^2 + y^2 - 2x + 4y - 4 = 0

OpenStudy (callisto):

Just change the constant term, k, of the equation to something else, as long as \(\sqrt{(-\frac{-2}{2})^2 + (-\frac{4}{2})^2-(k)} \ne 0 \) and \(\sqrt{(-\frac{-2}{2})^2 + (-\frac{4}{2})^2-(k)} \in \mathbb{R}\)